Find the velocity of an electron in the ground state of a hydrogen atom.
Answer #1


if electron is revalving around the nucleus in circular orbit

(1/2)*m*v^2=Ze^2/(4*pi*eo*r^2)

=====>

r=Ze^2/(4*pi*eo*m*v^2) .......(1)

and

from the Bohr's quantaization rule,

m*V*r=n*h/2*pi .............(2)

from (1) and (2)

v=z*e^2/(2eo*h*n)

for the hydrozen atom z=1 and for ground state n=1

v=1*e^2/(2eo*h*1)

=(1.6*10^-19)^2/(2*8.854*10^-12*6.626*10^-34)

=0.0218*10^8 m/sec

=2.182*10^6 m/sec